OpenStudy (anonymous):

solve the system of equations by eliminating the y variable: 9x+ 4y= -14 -2x+y=5

OpenStudy (anonymous):

Multiply all the terms in "-2x + y = 5" by -4 Giving 8x - 4y = -20, then add both equations 9x + 4y = -14 ------------ 17x = - 34 x = -2 and back solve for y 8x - 4(-2) = -20 8x =-20 - 8 x = -7/2

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

the wrong answer

OpenStudy (anonymous):

Plug the answer for x BACK into the equation. He plugged the x value into the y-variables.

OpenStudy (anonymous):

It should be: 8(-2) - 4y= -20 and continue to solve

OpenStudy (anonymous):

y=1

OpenStudy (anonymous):

Thanks Sam30317, that's a little more helpful than, " the wrong answer"

OpenStudy (anonymous):

and correct

OpenStudy (anonymous):

Lol!

OpenStudy (anonymous):

Do u understand how I got "1"?

jhonyy9 (jhonyy9):

9x+4y=-14 -2x+y=5 multiply by -4 and assum to first 9x+4y=-14 8x-4y=-20 ----------- 17x 0 =-34 x=-2 -2(-2)+y=5 4+y=5 y=1

jhonyy9 (jhonyy9):

hope this is right and understandably sure good luck bye

OpenStudy (anonymous):

When you plug in BOTH variables for either equation, you get -20 or -14